OpenStudy (anonymous):

Simplify the expression: 7^7 • 7 • 7^8 A. 7^56 B. 7^16 C. 7^15 D. 7^3

OpenStudy (anonymous):

OK, you have 7^7 so that is (7)(7)(7)(7)(7)(7)(7) * 7 * 7^8 which is 7 times itself 8 times and then solve

OpenStudy (anonymous):

it is faster if you solve on a calculator

OpenStudy (anonymous):

when you multiply exponents, the exponent rule is that you add the exponents

OpenStudy (anonymous):

using a calculator is cheating, jtryon!

OpenStudy (imstuck):

Are you familiar with, say, how to multiply the exponents when using a variable of x?\[x ^{5}\times x ^{3} \times x ^{2}\]It's the same as with variables. Keep the base of 7 and add the exponents.

OpenStudy (anonymous):

I know its faster but it's just wrong!

OpenStudy (anonymous):

Yeah, it is the exponent rule x^5x^6=x^5+6

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok...fine I don't get it!

OpenStudy (anonymous):

@adrynicoleb plz help!